# How Well Do Large Language Models Serve as End-to-End Secure Code Producers?

### v1

@@ Abstract.
The rapid advancement of large language models (LLMs) such as GPT-4 has revolutionized the landscape of software engineering, positioning these models at the core of modern development practices. As we anticipate these models to evolve into the primary and trustworthy tools used in software development, ensuring the security of the code they produce becomes paramount. How well can LLMs serve as end-to-end secure code producers? This paper presents a systematic investigation into LLMs’ inherent potential to generate code with fewer vulnerabilities. Specifically, We studied GPT-3.5 and GPT-4’s capability to identify and repair vulnerabilities in the code generated by four popular LLMs including themselves (GPT-3.5, GPT-4, Code Llama, and CodeGeeX2). By manually or automatically reviewing 4,900 pieces of code, our study reveals that: (1) large language models lack awareness of scenario-relevant security risks, which leads to the generation of over 75% vulnerable code on the SecurityEval benchmark; (2) LLMs such as GPT-3.5 and GPT-4 are unable to precisely identify vulnerabilities in the code they generated; (3) GPT-3.5 and GPT-4 can achieve 33.2%~59.6% success rates in repairing the insecure code produced by the 4 LLMs, but they both perform poorly when repairing self-produced code, indicating self-repair ”blind spots”. To address the limitation of a single round of repair, we developed a lightweight tool that prompts LLMs to construct safer source code through an iterative repair procedure based on the insights gained from our study. Experiments show that assisted by semantic analysis engines, our tool significantly improves the success rates of repair to 65.9%~85.5%.
@@ 1. Introduction
Transformer-based large language models111Note that there is currently no formal consensus on the minimum parameter scale for LLMs (Wang et al., 2024). This work focuses solely on larger language models with billion-scale parameters. have fundamentally reshaped software engineering in recent years. As LLM-powered programming assistants such as GitHub Copilot and CodeGeeX are widely adopted by IT companies and individual developers, traditional developer-centered software engineering (i.e., Software 1.0 and Software 2.0 where all source code is manually written) is rapidly evolving to LLM-centered software engineering (i.e., Software 3.0 where most source code is generated by AI). According to research by GitHub, its Copilot contributes to more than 46% of all source code that its 1.3 million users have developed (Dohmke, 2023).
@@
As more LLM-generated code is accepted by developers and thus becomes part of the software, its security is gaining increasing concern. Previous studies (Pearce et al., 2022a; Khoury et al., 2023; Nair et al., 2023) have revealed that although large language models are capable of generating functionally correct code, they are not free of security vulnerabilities, since their training sets contain real-world insecure code. Therefore, raw LLM-generated source code cannot be trusted to be deployed in security-sensitive scenarios.
@@
The increasingly significant role of LLMs in software engineering, coupled with disturbing vulnerabilities in the code they generate, compels us to explore methods for producing safer code with LLMs. Standard end-to-end software development practices, which include writing, reviewing, and refactoring code, result in significantly higher quality compared to focusing solely on code writing (Fowler, 2018). This inspires us to consider: How well do large language models function as comprehensive, end-to-end secure code producers? In the context of this research, end-to-end refers to a process wherein a code snippet is not only generated but also undergoes review and, if necessary, remediation of identified vulnerabilities before being presented to users. To achieve this level of competence, LLMs must demonstrate proficiency in three key areas: generating functionally correct code, conducting thorough self-reviews of their output, and effectively repairing any detected security flaws.
@@
Our investigation focuses specifically on evaluating these capabilities within the Python programming ecosystem. This choice is motivated by two primary factors: firstly, Python’s widespread adoption in web development, a domain where security considerations are paramount; and secondly, the relative scarcity of existing research examining the security implications of AI-generated Python code.
Hence, we seek answers to the following research questions (RQs):
RQ1
@@
How do LLMs perform when generating Python code in security-sensitive scenarios? This research question aims to address the lack of exclusive research on the security of LLM-generated Python code. Additionally, the identified insecure code in this RQ will be used in subsequent research questions.
RQ2
@@
How effective are LLMs in identifying LLM-generated code vulnerabilities? This research question seeks to investigate LLMs’ potential to self-review their generated code, as accurate identification of vulnerabilities is crucial for mitigating weaknesses in the source code.
RQ3
@@
How effective are LLMs in repairing LLM-generated code vulnerabilities? This RQ investigates LLMs’ capability to self-repair their generated code, which is the final and most crucial step in constructing secure source code.
RQ4
@@
How effective is an iterative strategy in improving LLMs’ repair capability? A single round of vulnerability detection and repair may not be sufficient to eliminate weaknesses in code. Therefore, we introduce an iterative strategy, allowing LLMs to repair the generated code through multiple attempts with feedback information provided
@@
Previous studies such as (Pearce et al., 2022a; Zhou et al., 2024a; Pearce et al., 2022b) have respectively examined the capability of large language models in generating secure code, detecting vulnerabilities in real-world code, and repairing them. However, these studies typically focus on a limited range of scenarios. Furthermore, there has been no comprehensive pipeline study integrating the processes of generation, detection, and repair to evaluate the potential of LLMs to produce secure code in an end-to-end manner.
@@
We present a systematic study evaluating the performance of 4 LLMs in generating secure Python code across 67 CWEs, offering a more extensive analysis than previous research as well as addressing the lack of relevant research focusing on Python. Overall, we found that the 4 tested LLMs generated over 75% vulnerable code on the SecurityEval benchmark.
To the best of our knowledge, we are the first to conduct a study on LLMs’ efficacy in judging and fixing insecure code generated by LLMs themselves, while previous work focuses primarily on manually written code in open-source projects. Our experiments revealed that GPT-3.5 and GPT-4 are unable to accurately identify weaknesses in LLM-generated code. While they achieve certain success rates in repairing the insecure code produced by the 4 LLMs (33.2% and 59.6% respectively), they perform poorer fixing self-produced code.
We designed an algorithm that prompts LLMs to produce much safer code through a simple feedback-driven iterative self-repair process. The algorithm achieves 65.9%~85.5% success rates of repair when using GPT-3.5 or GPT-4 as the code generator and repairer.
For the purpose of open science, we released the artifacts of this paper in our repository222https://github.com/jianian0318/LLMSecureCode.git.
@@ 2. Study Setting
2.1. Dataset
SecurityEval is a dataset dedicated to measuring the security of machine learning-based code generation techniques, proposed by Siddiq et al. (Siddiq and Santos, 2022) In its current version (sec, 2023), the dataset contains 121 code generation tasks in Python, all of which are exposed to the potential risk of a certain CWE. Overall, these tasks cover 69 CWEs. For reference, the dataset also provides an example vulnerable solution for each generation task.
As an example, under the CWE-089 directory of the dataset, there are two code generation tasks designed with potential security risk of CWE-89 (SQL Injection): CWE-089/author_1.py and CWE-089/codeql_1.py. Figure 1 shows CWE-089/author_1.py in the dataset, while Figure 2 is its example insecure solution vulnerable to SQL injection.
Overall, the dataset comprises 69 directories, each corresponding to a different CWE, with a total of 121 tasks under these directories.
We chose SecurityEval because (1) it was specially designed for evaluating the security of LLM-generated code, which is highly relevant to this work; (2) it covers the widest (by the time we conclude this research) range of 69 CWEs (121 scenarios), making our work more comprehensive than previous relevant ones (Pearce et al., 2022a; Khoury et al., 2023; Nair et al., 2023); (3) all of its code generation tasks are presented in Python, which matches our research goal.
@@ 2.2. Studied Large Language Models
Large language models capable of code processing tasks can be divided into 2 categories: general language models represented by the GPT family and specialized models that are specifically pre-trained on code (i.e., code language models) (Zhang et al., 2023). For better representation, we choose 2 popular models for our work from each category.
@@
(4) CodeGeeX2 is an open-source code language model with 6B parameters, developed jointly by Tsinghua University and Zhipu AI. It is based on the ChatGLM2 architecture and tailored for code-relevant tasks such as code generation, completion, interpretation, and cross-programming language translation (Cod, 2024; cod, 2024). CodeGeeX2 is chosen for our work because it is representative of a range of lightweight code language models.
@@ 2.3. Methodology
2.3.1. Methodology for RQ1~RQ3
Figure 3 shows the overall workflow of RQ1~RQ3, which goes in a pipeline fashion. The overall procedure aims to assess LLMs’ end-to-end capability to generate secure code.
In RQ1, we prompt GPT-3.5, GPT-4, Code Llama, and CodeGeeX2 to complete the 121 code generation tasks in SecurityEval. Results that do not meet basic functional requirements (with syntax errors or with obvious deviation from the intent of the prompt) will be regenerated until they do. We then review them to determine whether they are with the specific CWE vulnerabilities (explained in 2.3.3) and draw conclusions.
@@ 2.3.2. Algorithm for RQ4
To exploit the potential of LLMs in producing safer code, we design an iterative repair algorithm for RQ4, as shown in Figure 4. The algorithm takes code generation tasks as input. It initially utilizes a LLM to generate code according to given task descriptions. Subsequently, it employs an automated technique to scan the generated code for vulnerabilities. If a piece of code is free of weaknesses, it is output as secure code. Otherwise, it is returned to the LLM for repair, with CWE information provided. This scan-and-repair process is conducted iteratively until the code is secure or the predefined maximum number of iterations is reached.
It is worth noting that the ”automated scanning technique” in the algorithm is not predefined. Initially, we planned to use LLMs for vulnerability self-detection. However, as discovered in RQ2, LLMs are unable to provide reliable detection results. Alternatively, semantic analysis engines such as CodeQL and Bandit have been adopted.
@@ 2.3.3. Method of Code Review
Code snippets generated or repaired by LLMs need to undergo review to determine whether they are vulnerable to the specified CWEs. For instance, code generated for task CWE-089/author_1.py will be examined for the existence of CWE-89 (SQL Injection). We do not inspect the presence of other CWE vulnerabilities for (1) it is impossible for any method to encompass the entire spectrum of possible CWEs; (2) according to the design of SecurityEval, the generated code is most likely to exhibit the predefined vulnerabilities; and (3) focusing on predefined vulnerabilities is a common practice in previous studies such as (Pearce et al., 2022a; Siddiq and Santos, 2022; Pearce et al., 2022b). To ensure maximum accuracy, we perform 3 independent rounds of review for each piece of code:
(1) CodeQL Scan. CodeQL is an open-source semantic code analysis engine developed and maintained by GitHub, capable of detecting code vulnerabilities in various programming languages (cod, 2024). It is also the tool supporting GitHub Advanced Security features (git, 2024). CodeQL officially supports scanning (i.e., provides query scripts) for 29 CWEs and 1 CVE directly related to security in Python (que, 2024). Of these, 26 CWEs are among the 69 CWEs in SecurityEval, covering 67 out of 121 pieces of its code. We use CodeQL CLI v2.16.6 in our research. We choose CodeQL because of its status as an industry-leading engine for static code analysis and its widespread use in software security research.
(2) Bandit Scan. Bandit is a semantic code analysis tool specifically designed for Python by PyCQA (an organization for Python code standardization). Bandit is able to traverse code files, build abstract syntax trees, and report potential security vulnerabilities in CWE (ban, 2024). The version utilized in our experiment is Bandit 1.7.8. We chose Bandit for our research because it is tailored for Python, aligning closely with the focus of our work. While PyCQA has not officially announced Bandit’s coverage for CWE vulnerabilities, our practical experience reveals its capability to identify a range of vulnerabilities that may be overlooked by CodeQL.
(3) Manual Review. Given that CodeQL and Bandit do not comprehensively cover all the CWEs in SecurityEval, we conduct manual code review as a supplementary measure. In this process, the authors primarily refer to (1) the example insecure code provided by the SecurityEval dataset for each task, and (2) the CWE definitions available on the MITRE official website333https://cwe.mitre.org. Thereby, this approach ensures a thorough assessment of potential vulnerabilities beyond the scope of automated analysis tools.
We consider a piece of code to be vulnerable to its corresponding CWE if it has been identified as insecure in any of the three rounds of review, as described in Equation 1 where denotes whether the code is vulnerable.
@@ 2.4. Prompts and Parameters
This subsection describes the prompts (in zero-shot fashion) and parameters used in our research.
(1) Prompt for Code Generation. By referring to the OpenAI official website for guidance on code generation (OpenAI, 2023; pro, 2024), we construct the following prompt for code generation, where <Code> denotes a generation task in the SecurityEval dataset:
”You are a programming assistant, your task is to complete the following Python code according to comments faithfully: <Code>”
@@
(2) Prompt for Vulnerability Detection. We refer to relevant work on LLMs for vulnerability detection (Fu et al., 2023; Purba et al., 2023; Yin et al., 2024) to construct the following prompt used in RQ2:
”Your task is to determine whether the following Python code is vulnerable to CWE - <CWE Serial Number> (<CWE Definition>). Return True for vulnerable or False for invulnerable.”
@@
”You will be provided with a piece of Python code vulnerable to CWE - <CWE Serial Number> (<CWE Definition>). Your task is to generate the complete fixed code.”

### v2

# How Well Do Large Language Models Serve as End-to-End Secure Code Agents for Python?

@@ Abstract.
The rapid advancement of large language models (LLMs) such as GPT-4 has revolutionized the landscape of software engineering, positioning these models at the core of modern development practices. To fully realize their potential in producing secure source code autonomously, LLMs must not only generate code but also identify and repair vulnerabilities in their outputs, thereby improving security iteratively. Despite growing prominence, LLMs’ effectiveness in performing such end-to-end tasks remains unexplored. This paper bridges this gap by systematically investigating the capability of LLMs to generate source code, evaluate their own outputs for vulnerabilities, and apply necessary repairs to improve the security of their self-generated code.
Specifically, we studied the ability of GPT-3.5 and GPT-4 to identify and repair vulnerabilities in the code generated by four popular LLMs including themselves (GPT-3.5, GPT-4, Code Llama, and CodeGeeX2). By manually or automatically reviewing 4,900 pieces of code, our study reveals that: (1) LLMs generate over 75% vulnerable Python code in given scenarios; (2) LLMs such as GPT-3.5 and GPT-4 are unable to precisely identify vulnerabilities in the code they generated; (3) GPT-3.5 and GPT-4 can achieve 33.2%59.6% success rates in repairing the insecure code produced by the 4 LLMs, but they both perform poorly when repairing self-produced code, indicating self-repair “blind spots”. To address the limitation of a single round of repair, we developed a lightweight tool using LLMs as agents to construct safer source code through an iterative repair procedure based on the insights gained from our study. Experiments show that, assisted by semantic analysis engines, our tool significantly improves the success rates of repair to 65.9%85.5%.
| Related Work1 | Studied Language(s) | Automated Review | Manual Review | #Studied CWEs | #Studied LLM(s) |
| Pearce et al.(Pearce et al., 2022) | Python, C, Verilog | CodeQL | ✓ | 25 | 1 |
| Khoury et al.(Khoury et al., 2023) | C, C++, Python, HTML, Java | ✗ | ✓ | 14 | 1 |
| Asare et al.(Asare et al., 2023) | C, C++ | ✗ | ✓ | 6 | 1 |
| Sandova et al.(Sandoval et al., 2023) | C | ✗ | ✓ | 6 | 1 |
| Fu et al.(Fu et al., 2023a) | Python, JavaScript | CodeQL, Bandit, ESLint | ✗ | 16 | 3 |
| Liu et al.(Liu et al., 2024) | C, C++, Java | CodeQL | ✗ | 17 | 1 |
| Our work | Python | CodeQL, Bandit | ✓ | 69 | 4 |
1
Only studies which are based on code generated by LLMs and classify vulnerabilities according to CWEs are included in this comparison
@@ 1. INTRODUCTION
Transformer-based large language models have fundamentally reshaped software engineering in recent years. As LLM-powered programming assistants such as GitHub Copilot widely adopted by IT companies and individual developers, traditional developer-centered software engineering (i.e., Software 1.0 and Software 2.0 where all source code is manually written) is rapidly evolving to LLM-centered software engineering (i.e., Software 3.0 where most source code is generated by AI). According to research by GitHub, its Copilot contributes to more than 46% of all source code that its 1.3 million users have developed (Dohmke, 2023).
@@
As more LLM-generated code is accepted by developers and thus becomes part of the software, its security is gaining increasing concern. Previous studies (Pearce et al., 2022; Khoury et al., 2023; Nair et al., 2023; Asare et al., 2023; Sandoval et al., 2023; Fu et al., 2023a; Liu et al., 2024) have revealed that although large language models are capable of generating functionally correct code, they are not free of security vulnerabilities. Therefore, raw LLM-generated source code cannot be trusted to be deployed in security-sensitive scenarios.
@@
The increasingly significant role of LLMs in software engineering, coupled with disturbing vulnerabilities in the code they generate, compels us to explore methods for producing safer code with LLMs. Standard end-to-end software development practices, which include writing, reviewing, and refactoring code, result in significantly higher quality compared to focusing solely on code writing (Fowler, 2018). This inspires us to consider: How well do large language models function as comprehensive, end-to-end secure code agents? In the context of this research, end-to-end refers to a process wherein a code snippet is not only generated but also undergoes review and, if necessary, remediation of identified vulnerabilities before being presented to users. To achieve this level of competence, LLMs must demonstrate proficiency in three key areas: generating functionally correct code, conducting thorough self-reviews of their output, and effectively repairing any detected security flaws.
@@
Our investigation focuses specifically on evaluating these capabilities within the Python programming ecosystem. This choice is motivated by two primary factors: (1) Python’s widespread adoption in web development, a domain where security considerations are paramount; and (2) the scarcity of existing research focusing on the security implications of AI-generated Python code.
Hence, we seek answers to the following research questions:
@@
RQ1: How do LLMs perform when generating Python code in security-sensitive scenarios? This research question aims to address the lack of exclusive research on the security of LLM-generated Python code. Additionally, the identified insecure code in this RQ will be used in subsequent research questions.
@@
RQ2: How effective are LLMs in identifying LLM-generated code vulnerabilities? This research question seeks to investigate LLMs’ potential to self-review their generated code, as accurate identification of vulnerabilities is crucial for mitigating weaknesses in the source code.
@@
RQ3: How effective are LLMs in repairing LLM-generated code vulnerabilities? This RQ investigates LLMs’ capability to self-repair their generated code, which is the final and most crucial step in constructing secure source code.
@@
RQ4: How effective is an iterative strategy in improving LLMs’ repair capability? A single round of vulnerability detection and repair may not be sufficient to eliminate weaknesses in code. To address this, we introduce an iterative strategy that allows LLMs to repair the generated code through multiple attempts with feedback information provided. This research question synthesizes RQ1 through RQ3, serving as the focal point of our study.
@@
Previous studies such as (Pearce et al., 2022; Zhou et al., 2024a; Pearce et al., 2023) have examined, respectively, the ability of LLMs in generating secure code, detecting vulnerabilities in real-world code, or repairing them, similar to our effort in RQ1, RQ2, or RQ3. However, they suffer from limitations such as a narrow range of CWEs and restricted code review methods. In terms of RQ1, a comparison with representative work is detailed in Table 1. Furthermore, there has been no comprehensive pipeline study integrating the processes of generation, detection, and repair to evaluate the potential of LLMs to produce secure code in an end-to-end manner, a gap that our RQ4 aims to address.
@@
(1) We present a systematic study evaluating the performance of 4 LLMs in generating secure Python code across 67 CWEs, offering a more extensive analysis than previous research as well as addressing the lack of relevant research focusing on Python. Overall, we found that the 4 tested LLMs generated over 75% vulnerable code on the SecurityEval benchmark. Meanwhile, we made several novel discoveries such as self-repair ”blind spots” of LLMs.
(2) To the best of our knowledge, we are the first to conduct a study on LLMs’ efficacy in judging and fixing insecure code generated by LLMs themselves, while previous work focuses primarily on manually written code in open-source projects. Our experiments revealed that GPT-3.5 and GPT-4 are unable to accurately identify weaknesses in LLM-generated code. While they achieve certain success rates in repairing the insecure code produced by the 4 LLMs (33.2% and 59.6% respectively), they perform poorer fixing self-produced code.
(3) We designed a tool using LLMs as agents to produce much safer code through a simple feedback-driven iterative self-repair process. The tool achieves 65.9%85.5% success rates of repair when using GPT-3.5 or GPT-4 as the code generator and repairer.
For the purpose of open science, we released the artifacts of this paper in our repository111https://github.com/jianian0318/LLMSecureCode.
@@ 2. STUDY SETTING
2.1. Studied Dataset
SecurityEval is a dataset dedicated to measuring the security of machine learning-based code generation techniques(Siddiq and Santos, 2022). In its current version, the dataset contains 121 code completion tasks in Python, all of which are exposed to the potential risk of a certain CWE. Overall, these tasks cover 69 CWEs. The dataset also contains an example vulnerable solution for each generation task. Fig. 1 depicts the structure of SecurityEval.
We chose SecurityEval because (1) it was specially designed for evaluating the security of LLM-generated code, which is highly relevant to this work; (2) it covers the widest (by the time we conclude this research) range of 69 CWEs (121 scenarios), making our work more comprehensive than previous relevant ones (Pearce et al., 2022; Khoury et al., 2023; Nair et al., 2023); (3) all of its code generation tasks are presented in Python, which matches our research goal.
@@ 2.2. Studied Large Language Models
(4) CodeGeeX2 is an open-source code language model with 6B parameters, developed jointly by Tsinghua University and Zhipu AI. It is based on the ChatGLM2 architecture and tailored for code-relevant tasks such as code generation, completion, interpretation, and cross-programming language translation (Cod, 2024)(cod, 2024). CodeGeeX2 was selected for our study because of its strong performance among lightweight LLMs for code (Zheng et al., 2023).
@@ 2.3. Methodology
2.3.1. Methodology for RQ1RQ3
Fig. 2 shows the overall workflow of RQ1RQ3, which goes in a pipeline fashion. The overall procedure aims to assess LLMs’ end-to-end capability to generate secure code.
In RQ1, we prompt GPT-3.5, GPT-4, Code Llama, and CodeGeeX2 to complete the 121 code generation tasks in SecurityEval. Results with syntax errors will be re-generated until free of errors. We then review them to determine whether they are with the specific CWE vulnerabilities (explained in 2.3.3) and draw conclusions.
@@ 2.3.2. Algorithm for RQ4
To exploit the potential of LLMs in producing safer code, we design an iterative repair algorithm for RQ4, as shown in Algorithm 1. Taking code generation tasks () as input, it initially utilizes a LLM to generate code (LLM_gen()) according to given task descriptions. Subsequently, it employs an automatic tool (scanner(), CodeQL and Bandit are used in RQ4) to scan the generated code for vulnerabilities. If a piece of code is free of weaknesses, it is output as secure code. Otherwise, it is returned to the LLM for repair (LLM_repair()), with CWE information provided (). This scan-and-repair process is conducted iteratively until the code is secure or the predefined maximum number of iterations () is reached.
It is worth noting that the “scanner” in the algorithm is not predefined. Initially, we planned to use LLMs for vulnerability self-detection. However, as discovered in RQ2, LLMs are unable to provide reliable detection results. Alternatively, semantic analysis engines such as CodeQL and Bandit have been adopted.
@@ 2.3.3. Method of Code Review
Code snippets generated or repaired by LLMs need to undergo review to determine whether they are vulnerable to the specified CWEs. For instance, code generated for task CWE-089/author_1.py will be examined for the existence of CWE-89 (SQL Injection). We do not inspect the presence of other CWE vulnerabilities for (1) it is impossible for any method to encompass the entire spectrum of possible CWEs; (2) according to the design of SecurityEval, the generated code is most likely to exhibit the predefined vulnerabilities; and (3) focusing on predefined vulnerabilities is a common practice in previous studies such as (Pearce et al., 2022; Siddiq and Santos, 2022; Pearce et al., 2023). To ensure maximum accuracy, we perform 3 independent rounds of review for each piece of code:
(1) CodeQL is an open-source semantic code analysis engine developed and maintained by GitHub, capable of detecting code vulnerabilities in various programming languages (cod, 2024). It is also the tool supporting GitHub Advanced Security features (git, 2024). CodeQL officially supports scanning (i.e., provides query scripts) for 29 CWEs and 1 CVE directly related to security in Python (que, 2024). Of these, 26 CWEs are among the 69 CWEs in SecurityEval, covering 67 out of 121 pieces of its code. We choose CodeQL because of its status as an industry-leading engine for static code analysis and its widespread use in software security research. The version we used is CodeQL CLI v2.16.6.
(2) Bandit is a semantic code analysis tool specifically designed for Python by PyCQA (an organization for Python code standardization). Bandit is able to traverse code files, build abstract syntax trees, and report potential security vulnerabilities in CWE (ban, 2024). We chose Bandit for our research because it is tailored for Python, aligning closely with the focus of our work. The version utilized in our experiment is Bandit 1.7.8.
(3) Manual Review. Since CodeQL and Bandit do not comprehensively cover all CWEs in the SecurityEval dataset, 4 reviewers were tasked with conducting independent manual code reviews as a supplementary measure. During this process, the reviewers primarily referred to the example insecure code provided by SecurityEval for each task, as well as the CWE definitions available on the official MITRE website222https://cwe.mitre.org. In cases of disagreement, the final decision was made by an expert in Software Engineering following a group discussion with the reviewers. This approach ensures a thorough evaluation of potential vulnerabilities beyond the scope of automated analysis tools.
We consider a piece of code to be vulnerable to its corresponding CWE if it has been identified as insecure in any of the three rounds of review, as described in (1) where denotes whether the code is vulnerable.
@@ 2.4. Prompts and Parameters
This subsection describes the prompts (in zero-shot fashion) and parameters used in our research. Due to the large number of cases and the significant effort required for manual review, we only focused on the performance of the LLMs under the following set of prompts:
(1) Prompt for Code Generation. By referring to the OpenAI official website for guidance (OpenAI, 2023; pro, 2024), we construct the following prompt for code generation, where <Code> denotes a generation task in the SecurityEval dataset:
“You are a programming assistant, your task is to complete the following Python code according to comments faithfully: <Code>”
@@
(2) Prompt for Vulnerability Detection.
We refer to relevant work on LLMs for vulnerability detection (Fu et al., 2023b; Purba et al., 2023; Yin et al., 2024) to construct the following prompt used in RQ2, in which <SN> and <Def> respectively denote the serial number and the definition of the CWE:
“Your task is to determine whether the following Python code is vulnerable to CWE-<SN>(<Def>). Return True for vulnerable or False for invulnerable.”
@@
“You will be provided with a piece of Python code vulnerable to CWE-<SN>(<Def>). Your task is to generate the complete fixed code.”